# Identifying Reliable Predictions in Detection Transformers

### v1

@@ Abstract
DEtection TRansformer (DETR) has emerged as a promising architecture for object detection, offering an end-to-end prediction pipeline. In practice, however, DETR generates hundreds of predictions that far outnumber the actual number of objects present in an image. This raises the question: can we trust and use all of these predictions? Addressing this concern, we present empirical evidence highlighting how different predictions within the same image play distinct roles, resulting in varying reliability levels across those predictions. More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated. Based on these insights, we demonstrate identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
@@
Building on this viewpoint, we first tackle the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error. Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized. In response, we present Object-level Calibration Error (OCE), which is capable of assessing the calibration quality both across different models and among various configurations within a specific model. As a final contribution, we introduce a post hoc Uncertainty Quantification (UQ) framework that predicts the accuracy of the model on a per-image basis. By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, the framework assesses the reliability of the DETR model for each test image. The code is available at https://github.com/azizanlab/reliable-detr.
(a) Low confidence with random boxes
(b) Calibrated confidence with similar boxes
(c) Uncalibrated confidence with similar boxes
@@
Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with one being well-calibrated. This figure illustrates how DETR can handle the remaining predictions. In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with similar boxes, or (c) poorly calibrated confidence with similar boxes. Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
1 Introduction
@@
Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis. Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [8, 23, 22, 9] to identify and locate objects within images. However, the recent introduction of DEtection TRansformer (DETR) [1] has revolutionized the field by offering an end-to-end prediction pipeline where the model predicts a set of bounding boxes and class probabilities.
The core innovation of DETR lies in the use of a Transformer encoder-decoder architecture, enabling the model to generate predictions in an end-to-end manner and enhancing scalability. This paradigm shift has led to the exploration of various DETR variants—such as Deformable-DETR [32] and DINO [31]—positioning them as potential foundation models for object detection tasks. Despite these advancements, the inner workings of how these predictions are generated and interact within the Transformer decoder layers remain under-explored.
@@
Likewise, there is an ongoing debate within the community about whether DETR truly provides an end-to-end prediction pipeline. The central concern is whether all the hundreds of predictions generated by DETR can be trusted and used. By means of an answer, practitioners often employ heuristic approaches to achieve high precision, such as by setting a user-defined threshold to retain only a small subset of high-confidence outputs, as seen in the official demo. Previous studies exploring model reliability, such as [10, 15, 16, 21, 17], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set. On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR and 300 out of 900 for DINO) based on confidence scores during post-processing to achieve a high average precision (AP) score. However, the significance of selecting a subset of the predictions as well as the choice of a suitable configuration for doing so remain under-explored. We provide a more in-depth discussion about related works in Appendix A.
This paper examines these issues by addressing the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability? (RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set? and (RQ3) How can we accurately assess DETR’s reliability?
@@
Our primary findings reveal that predictions from DETR within the same image are interdependent, leading to varying levels of reliability. Since DETR is trained using gradients from optimally matched predictions, it has flexibility in handling the remaining predictions. In principle, DETR could assign them low confidence with random bounding boxes (Figure 1(a)). Alternatively, it could also generate similar bounding boxes with equally calibrated confidence (Figure 1(b)). Otherwise, it may assign poorly calibrated confidence scores (Figure 1(c)). Our analysis indicates that DETR mostly follows the third scheme; we refer to these well-calibrated predictions as positive predictions and the others as negative predictions.
Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and how this finding can be leveraged into an uncertainty quantification (UQ) framework for DETR. More specifically, our main contributions are:
@@
  • We offer both qualitative and quantitative insights into how positive and negative predictions influence the model’s object-level calibration. Our analysis shows that negative predictions are poorly calibrated. Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use.
  • We raise concerns about existing performance and calibration metrics—including AP [25, 6, 14], detection expected calibration error (D-ECE) [10], and Localisation-aware expected calibration error (LA-ECE) [19, 11]—by noting their inadequacy for identifying the reliable subset of predictions. For instance, AP favors retaining more predictions, while ECE prefers a smaller high-confidence set. Consequently, the subset of predictions that maximizes performance for each metric differs significantly from the optimal positive set.
@@
  • To address this, we introduce a new calibration metric, object-level calibration error (OCE), which calculates calibration error along ground truth objects rather than predictions. We demonstrate that OCE is capable of both ranking different models by their calibration qualities and identifying the reliable subset of predictions to use.
@@
  • Based on our findings, we provide a practical implication: we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric. We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios.
@@
  • We present a comparative analysis of methods for identifying positive predictions in DETR, from the UQ perspective. Our results show that thresholding (on predictive confidence score) is a more suitable method than widely used techniques such as top-$k$ or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
2 Problem Statement
2.1 Detection Transformers (DETRs)
@@
We consider a test image $\bm{x}$ and denote $\mathcal{T}_{\bm{x}}$ as the set of ground truth objects present in the image. Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{T}}_{\theta}(\bm{x})$. Each prediction $t\in\hat{\mathcal{T}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
@@
The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions. In addition to the features extracted by the encoder, the decoder’s input consists of $M$ (typically several hundred) learnable embeddings, also known as object queries. Each decoder layer is composed of a self-attention module among object queries and a cross-attention module between each object query and the features. After processing the queries through several decoder layers, the model produces the $M$ final representation vectors that are converted into bounding boxes and class labels via a shared feedforward network, $f_{\phi}$. Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set. We refer to Figure 4 in the Appendix for an illustration.

### v2

# Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty

@@ Abstract
DEtection TRansformer (DETR) has emerged as a promising architecture for object detection, offering an end-to-end prediction pipeline. In practice, however, DETR generates hundreds of predictions that far outnumber the actual number of objects present in an image. This raises the question: can we trust and use all of these predictions? Addressing this concern, we present empirical evidence highlighting how different predictions within the same image play distinct roles, resulting in varying reliability levels across those predictions. More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated. Based on these insights, we demonstrate that identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
@@
Building on this viewpoint, we first address the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error. Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized. In response, we present Object-level Calibration Error (OCE), which assesses the calibration quality more effectively and is suitable for both ranking different models and identifying the most reliable predictions within a specific model. As a final contribution, we introduce a post hoc uncertainty quantification (UQ) framework that predicts the accuracy of the model on a per-image basis. By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, our framework assesses the reliability of the DETR model for each test image.
(a) Random boxes. Low Confidence.
(b) Accurate boxes. Calibrated.
(c) Accurate boxes. Uncalibrated.
@@
Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated. This figure illustrates how DETR can handle the remaining predictions (i.e., red and gray boxes). In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with accurate prediction for the bounding boxes, or (c) poorly calibrated confidence with accurate boxes. Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
I Introduction
@@
Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis. Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [1, 2, 3, 4, 5, 6] to identify and locate objects within images. More recently, DEtection TRansformer (DETR) [7] has revolutionized the field by utilizing a Transformer encoder-decoder architecture to offer a scalable end-to-end prediction pipeline, where the model predicts a set of bounding boxes and class probabilities. This paradigm shift has led to the exploration of various DETR variants, positioning them as potential foundation models for object detection tasks. While notable progress has been made, the reliability of these predictions remains insufficiently investigated.
The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used. For example, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions. As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler. Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image. However, owing to DETR’s unique set-prediction mechanism, quantifying its image-level reliability is not straightforward and remains largely underexplored.
@@
Moreover, there is an ongoing debate within the community about whether DETR can be considered an entirely end-to-end solution, free from any post-processing requirements in practice. In particular, DETR outputs a fixed number of predictions, typically in the hundreds; consequently, the central concern is which predictions can be trusted and used. Practitioners often employ heuristic approaches to determine this subset, such as by setting a user-defined threshold to retain only a small subset of high-confidence (e.g., $>0.7$) outputs, as seen in the official demo. Similarly, previous studies exploring model reliability, such as [8, 9, 10, 11, 12, 13], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set. On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence score. Nonetheless, how different subset selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remain under-explored, and their significance has yet to be fully recognized.
@@
Our primary findings reveal that DETR predictions within the same image are interdependent, leading to significantly varying levels of reliability. Since DETR is trained using gradients derived exclusively from optimally matched predictions (i.e., the predictions most closely corresponding to each annotated object), it will output at least one well-calibrated prediction for each perceived object while having flexibility in handling the remaining unmatched predictions. Throughout the paper, we refer to those matched predictions as positive predictions and the others as negative predictions. In principle, to avoid an inaccurate match, DETR could assign those unmatched predictions low confidence with random bounding boxes (Figure 1a). Alternatively, it could generate similarly accurate bounding boxes with (nearly) equally calibrated confidence (Figure 1b). Otherwise, it may assign poorly calibrated confidence scores (Figure 1c). Our analysis indicates that a prediction generated by DETR mostly follows the third scheme.
Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and addresses the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability? (RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set? and (RQ3) How can we accurately assess DETR’s image-level reliability? In addressing these questions, our specific contributions include:
  1. 1.
@@
    We provide both qualitative and quantitative insights into how positive and negative predictions influence the model’s reliability. Our analysis reveals that, unlike positives, negative predictions are often poorly calibrated and therefore should be handled separately. Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use (Section IV).
  2. 2.
@@
    Identifying ground-truth positives and negatives ideally requires a matching process based on ground-truth annotations, which are unavailable at test time. Consequently, an alternative framework for identifying positive predictions is needed. However, we show that existing performance and calibration metrics—such as average precision [16, 17, 18] and expected calibration error [8, 19, 20]—are inadequate for this purpose. To address this, we introduce a new calibration metric, object-level calibration error (OCE), which assesses calibration error along ground-truth objects rather than predictions. We demonstrate that OCE is more suitable for both ranking different models by their calibration qualities and identifying the positive predictions (Section V).
  3. 3.
@@
    Based on our findings, we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric. We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios (Section VI).
@@
In addition to addressing the three primary research questions, we present a comparative analysis of standard post-processing methods for identifying positive predictions in DETR from an uncertainty quantification (UQ) perspective (Section VII). Our results indicate that thresholding on the predictive confidence score is more effective than widely used techniques such as top-$k$ selection or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
II Related Work
Calibration. Model calibration refers to how well a model’s predicted confidence scores align with the actual likelihood of those predictions being correct. In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy. For instance, if a model predicts an event with 80% confidence, it should be correct around 80% of the time for that event to be considered well-calibrated. Achieving good calibration is crucial because it improves trust in the model’s predictions, guides more informed decision-making, and enables better risk assessment in real-world applications.
To evaluate the alignment, they measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin. However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes. Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction. To address these challenges, D-ECE [8] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$). On the other hand, LaECE [21] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
However, as we demonstrate, DETR exhibits poor calibration on negative predictions, regardless of its actual reliability. In practice, users often focus only on positive samples. Yet, when model calibration is evaluated across the entire set of predictions, including the negatives, the resulting evaluation becomes highly biased and therefore unreliable. To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$. As detailed in later sections, a threshold of $0.3$ approximates the optimal positive predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable. However, the optimal threshold is not guaranteed to be always $0.3$ for other models like UP-DETR, which introduces potential risks. In response, our paper proposes OCE, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
Uncertainty Quantification. Several studies focus on out-of-distribution (OOD) identification in object detection models. For example, [22] propose a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions. [23] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data. Similarly, [19] employ an auxiliary detection model capable of expressing its confidence. Other works, including [24] and [25], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
To the best of our knowledge, the aforementioned existing UQ techniques primarily focus on prediction-level analysis. Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection models. In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs. Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
III Preliminaries
Figure 2: A diagram of the DETR architecture. An input image is first processed through a CNN backbone to generate a 2D feature representation. This representation is then passed to the Transformer encoder, which extracts feature vectors. These feature vectors are sent to the decoder, which receives $M$ learned object queries together. The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
III-A Detection Transformer (DETR)
@@
We consider a test image $\bm{x}$ and denote $\mathcal{D}_{\bm{x}}$ as the set of ground truth objects present in the image. Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{D}}_{\theta}(\bm{x})$. Each prediction $d\in\hat{\mathcal{D}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
@@
The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions. In addition to the features extracted by the encoder, the decoder’s input consists of $M$ (typically several hundred) learnable embeddings, also known as object queries. Each decoder layer is composed of a self-attention module among object queries and a cross-attention module between each object query and the features. After processing the queries through several decoder layers, the model produces the $M$ final representation vectors that are converted into bounding boxes and class labels via a shared feedforward network, $f_{\phi}$. Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set. We refer to Figure 2 for an illustration.